What is the frequency of light in a vaccum that has a wavelength of 7.14x10 4m
1
Expert's answer
2020-04-23T12:50:16-0400
"f=\frac{c}{\lambda}=\frac{3\cdot10^8}{7.14\cdot10^4}=4200\ Hz=4.2\ kHz"